OpenStudy (anonymous):

newtons second law of motion if the force on an object is 20N and its mass 5 g, what is its acceleration?

OpenStudy (anonymous):

We know that newton's second law is expressed as\[F = m \cdot a\] Simply substitute in the known values (F = 20N, m = 0.005kg) and solve for a.

OpenStudy (anonymous):

F=m*a 20=.005*a a=4000ms^2